Whirly Corporation’s contribution format income statement for the most recent month is shown below: Total Per Unit Sales (7,800 units) $ 249,600 $ 32.00 Variable expenses 148,200 19.00 Contribution margin 101,400 $ 13.00 Fixed expenses 54,600 Net operating income $ 46,800 Required: (Consider each case independently): 1. What would be the revised net operating income per month if the sales volume increases by 30 units? 2. What would be the revised net operating income per month if the sales volume decreases by 30 units? 3. What would be the revised net operating income per month if the sales volume is 6,800 units?
Answer #1
Case 1 Case 2 Case 3
Unit sales (7800+30)=7830 (7800-30)=7770 6800
Sales (7830*32)=250560 (7770*32)=248640 (6800*32)=217600
Variable expenses (7830*19)=148770 (7770*19)=147630 (6800*19)=129200
Contribution margin 101790 101,010 88400
Fixed expenses 54600 54600 54600
Revised net operating income $47190 $46410 $33800

NOTE:Total fixed costs and variable cost per unit do not change with change in units